From ObsVermes
Jump to: navigation, search

The news of the years 2008, 2009 and 2010 are stored in the archive.

Door open again (2016-10-16)

After Lukas wrote the correct udev rule the door opened.
Focus run at low temperatures.

After a period of inactivity at Obs. Vermes we could manage to get all the components up and running again.

One of the reasons was my winter over stay at Dome Concordia in Antarctica as a member of DC 11. I took care of the infrared telescope IRAIT/ITM and of the camera AMICA, wrote a new RTS2 driver for the mount and acquired data for a pointig model and of course I focused IRAIT/IT with rts2saf. Living and working at an altitude of 3227 m or better under a pressure of 630 hPa, temperatures as low as -81.4 degC and a longer period of almost complete darkness was tough. Fortunately the DC11 crew teamed up and made it easy to endure the nine months of complete isolation. Living on almost nothing but a flat ice shield may appear boring but to me the opposite was the case, there were so many things to discover.

All had begun at the conference Autonomous Observatory Software RTS2 Lijiang, China, 2014 where I met again Jean-Marc, the PI of IRAIT/ITM, from Aosta Valley. After a short conversation and after talking with my wife it was clear that I'd go to Antarctica.

rts2af@Bootes-2,Málaga (2013-07-03)

Offline analysis of a focus run taken on 2013-07-03.

While the weather in Switzerland is still bad, I took the chance to visit Bootes-2 near Málaga to install and test rts2af package. Last night I took several samples. Adjustments to SExtractor parameters were necessary in order to receive enough objects on each image. Done that, rts2af reproduced the manually chosen best focus position (1645 ticks) as 1642 ticks. The FWHM was 7.2 pixels.

Polar alignment measurement with rts2pa (2012-09-28)

Last image taken at 2012-09-28T19:25:16.014, conversion to PNG done with fitspng

On 2012-09-17 the Internet connection was interrupted since did not reboot properly. After manual boot the observatory resumed operation.

In the evening the script took five images and the result of the polar alignment calculation is

KingA dx=-5.064 dy=-17.1044996475 arcsec
KingA dtau=0.135778986508 arcsec RA= 19.7235488889 HA=2.93587219088
KingA HA=350.333610323 deg lambda=131.378103146 arcsec A=-32.5488058658, k=129.5

The HA axis is tilted about 2 minutes to the East above the north celestial pole (NCP). Within 11 years, the date of the last measurement was 2001-09-30, the HA axis moved about by 60-80 arcsec mostly to the South.

The original images are available here.

Downs, ups, rts2af, rts2pa, (2012-08-05)

The inverter of the door motor died while the door was open. That required a taxi ride on 2012-07-09 at 23:30. This device has been replaced and since 2012-08-03 the observatory is fully operational again.

Since early June I support Baker Observatory in getting the rts2af autofocuser package up and running. Main of the enhancement tasks were better test features. It is now possible to test the whole chain with existing images] together with a running RTS2 system. A new feature is the fall back calculation for best focus based on weighted mean of FWHM and flux.

Polar alignment can now be measured with the help of rts2pa package.

For both packages installation, testing instructions as well as man pages describing the general operation and configuration are available.

Obs. Vermes operational (2012-07-07)

After a long repair of FLI PDF and update of the computer infrastructure Obs. Vermes is now operational again. (2012-04-15)

During the past weeks the gateway reported that its disk was dying. The Alix host was reconfigured to run OpenSuSE 12.1 and strongSwan replaced Openswan. It turned out that strongSwan itself is easier to configure and to debug than Openswan, but to resolve the interplay with openSuSE's SuSEfirewall2 was still an undocumented adventure.

Obs. Vermes down (2012-03-24)

Again the FLI PDF focuser ceased its operation. Fortunately(?) this time the motors don't have enough torque to drill out the threads.

Obs. Vermes up (2012-03-17)

The new FLI PDF focuser.

After inspection it turned out that the steel screws destroyed the threads and a complete replacement of the focuser was necessary.

Obs. Vermes down (2012-02-08)

The two red markers show clearly the jammed screws, really screwed up :-((.

The Obs. Vermes is down due to a jammed FLI PDF focuser. This is the second time that this product caused a failure.

Extended temperature model to -16.2C (2012-02-04)

Extended range from -2C down to -16.2C.

Due to the exceptionally cold weather (outdoor temperature -22C) the focuser temperature model has now a data range from -16.2C to 25C. The whole observatory operated all night without any malfunction.

Obs. Vermes exiled (2012-01-29)

View in direction of Tromsø at 2012-01-24 23:36:05 CET from our hut Aslund.

As announced on 2012-01-24 evening we saw extra ordinary northern lights (aurora borealis,G1-class geomagnetic storm) in Norway near Jøvik (Longitude: 19.722473 E, Latitude: 69.544480) having a clear sky.

Typical settings for NIKON D70s:

 ExifTool Version Number: 8.40 
 File Name              : DSC_0019.JPG
 Camera Model Name      : NIKON D70s
 Orientation            : Horizontal (normal)
 Exposure Time          : 15
 F Number               : 4.5
 Date/Time Original     : 2012:01:24 23:36:05
 Focal Length           : 18.0 mm           
 ISO                    : 800          
 Quality                : Basic   
 White Balance          : Preset 
 Shooting Mode          : IR Control 
 Sensor Pixel Size      : 7.8 x 7.8 um   
 Lens ID                : AF-S DX Zoom-Nikkor 18-70mm f/3.5-4.5G IF-ED
 Create Date            : 2012:01:24 23:36:05.80
 Circle Of Confusion    : 0.020 mm
 Field Of View          : 67.3 deg (13.31 m)

Storm front Joachim (2011-12-17)

Storm front Joachim interrupted the power supply for about 12 hours. UPS went low and NUT shut RTS2 and all computers down. Resuming operation revealed that there was no IP connection to the equipment on the dome. The power supply of Linksys WET54G died and we replaced it.

Obs. Vermes up again (2011-11-26)

The ssd650v frequency inverter says rdy.

Today we installed the new frequency inverter and since the weather was good started immediately the observations. Out of the box the FWHM was 4.2 pixels

2011-11-26 18:56:17,085 INFO FWHM: 3115.0, 4.2803565365, 589, X, /scratch//images//que/20111126/20111126175506-148-RA.fits

for a white light image and after the first focus run

2011-11-26T19:22:52.289 CET EXEC 4 rts2af_acquire: focus line: FOCUS: 3426.943141, FWHM: 2.393442, TEMPERATURE: 4.94444417953C, OBJECTS: 64 DATAPOINTS: 11

the focus was found at 3427 and the FWHM was 2.4 pixels. Subsequent white light images have a FWHM varying between 1.8 ... 4.2 pixels (indicating room for improvements).

Obs. Vermes down again (2011-11-11)

The new ssd650v frequency inverter.

Since 2011-10-22 Obs. Vermes is out of order because of a total failiure of dome azimuth frequency inverter. The new inverter arrived today.

Obs. Vermes resumed operation (2011-10-01)

A focus run "out of the box".

After two months Obs. Vermes started its normal operation. It's pitty we had so many good nights. Update 2011-10-02: and it got the focus again.

Obs. Vermes down (2011-07-31)

Finally the first frequency inverter arrived in bad condition

The ssd650v frequency inverter went off line last Thuesday (2011-07-26) and the site is not reachable since last Saturday. A visit this week will be necessary.
Update 2011-08-07: had a erroneous file system preventing normal boot process. The inverter SSD650v communicates not reliably over RS 232 anymore. It will be replaced.

Update: The first frequency inverter arrived on last Friday (2011-09-23) in bad condition (see image) and the second on Monday (2011-09-25). But it seems that the software changed or it does not accept the command to set the setpoint (247).

Mean of the fitted functions (2011-07-23)

Explanation see text.

In order to further clarify the open issue mentioned in the next paragraph the analysis has been extended and the mean of the parameters of the fitted functions have been determined. On the graph at the lower left the green lines indicate the individual and the red curve the mean of the fitted functions. In the region of the minimum (+/-100 ticks) the FWHM variation is about .2 pixels and hence a slightly off focus position can not explain the growing differences below.

In order to compare the fitted polynomials the minimum has been determined first. Then the FOC_POS_reduced has been calculated as (FOC_POS- FOC_POS_MIN). p0r, p1r, p2r are the parameters of the fitted even 4th degree polynomial.

Up to now all focus runs were done in the white light. To reduce the errors only objects near the center with a radius less than 800 1x1 pixels have been used for this analysis. To find the parameters of the fitted function it might be better to use a (narrow) filter in the optical train. Comparing filter - filter offsets should yield higher precision, but then still the offset to the white light needs careful attention.

More experience (2011-07-10)

A more practical approach: FOC_POS(T).

The problems with setting the focuser position have been sorted out and only in rare cases the FLI PDF does not reach the target position. Most of the time the focuser was operating in relative mode. Relative mode means that the FWHM of an image is analyzed and used to decide to schedule a run if it is above threshold. The normal procedure should be that a focuser run occurs in the maximum once per night (e.g. at the beginning) and then temperature compensation takes over the control. The night of 2011-07-05 was not particularly clear so the explanation for the increasing FWHM could indeed be denser and denser clouds. The sudden FWHM drop at 03:12:07 and the more or less constant fitted FWHM stand in contrast to this idea but it can not be decided without the narrow angle cloud sensor.

2011-07-05 00:53:22,444 INFO FWHM:4.31583333333
2011-07-05T01:01:17.359 rts2af_acquire: FOCUS: 3680.055732, FWHM: 2.358561, TEMPERATURE: 16.2777786255C
2011-07-05 01:03:16,463 INFO FWHM:2.36775862069
2011-07-05 01:09:07,437 INFO FWHM:4.37724137931
2011-07-05 01:14:13,156 INFO FWHM:3.75842105263
2011-07-05 01:19:19,785 INFO FWHM:3.81375
2011-07-05 01:24:23,093 INFO FWHM:3.8525
2011-07-05 02:04:33,777 INFO FWHM:5.77862068966
2011-07-05 02:06:22,956 INFO FWHM:5.51906666667
2011-07-05 02:11:20,868 INFO FWHM:5.85329113924
2011-07-05T02:16:47.523 rts2af_acquire: FOCUS: 3702.626771, FWHM: 2.490576, TEMPERATURE: 16.2777786255C
2011-07-05 02:17:31,968 INFO FWHM:5.68388349515
2011-07-05 02:18:55,127 INFO FWHM:3.075625
2011-07-05 02:30:45,343 INFO FWHM:4.22466666667
2011-07-05 02:39:29,658 INFO FWHM:3.664
2011-07-05 02:50:07,742 INFO FWHM:4.89428571429
2011-07-05 03:01:07,111 INFO FWHM:5.215
2011-07-05T03:07:03.303 rts2af_acquire: FOCUS: 3603.294323, FWHM: 2.500838, TEMPERATURE: 14.8333330154C
2011-07-05 03:12:07,213 INFO FWHM:1.94508571429
2011-07-05 03:25:24,375 INFO FWHM:3.63823529412
2011-07-05 03:36:20,648 INFO FWHM:3.33786885246

The hot end (2011-06-29)

Extending the range between 20 and 27C.

We had a few days with a maximum temperature above 30 degC and during the night it did not drop below 16 degC. This was a good occasion to extend the range towards the "hot" end. For the graph on the right side the selection criteria were slightly modified. It turned out that a fitted FWHM below 2.3 pixels contribute overly to the tails of the residuals.

Looking at the data the absolute resolution of the the focuser is about 100 ticks or 2 degC over a period of more than six months.

At the RTS2 repository is a temperature compensating version flitc.cpp available for the FLI PDF. It is not yet stable since there are still unresolved issues concerning the setting of the focuser position (2011-07-10: most bugs in rts2af_acquire removed).

Temperature compensation for FLI focuser (2011-06-18)

Today the temperature compensating RTS2 FLI focuser driver flitc.cpp has been released to RTS2 repository.

Data since 2010-12-01, left and middle temperature from Davis weather station (out-, indoor), right IRS sensor temperature from AAG cloud device.

Temperature model for AP Astrofire F/7 (2011-06-03)

Retrieving the data from the local meteo database yielded a significant improvement of the temperture dependency of the focuser position. The constraints

  • more than 10 objects
  • a chi2 of less than 5.
  • a fitted FWHM of less than 4.0 pixel (1.41 arcsec/pixel)

are moderate. The analysis is still in progress.

96 focuser runs taken through the past six months.

Initial temperature model for AP Astrofire F/7 (2011-05-29)

Due to the large FWHM variation as a function of ambient temperature frequent focus run were observed. To compensate the temperature variation the primary aim is to establish a model for the AP Astrofire F/7 refractor which can be used by the focuser to predict a reasonable position.

The developed software infrastructure uses the existing rts2af package. Two new scripts and have been released to the repository. The first scans a given base directory for focus runs of a given filter (filter X, that is the clear optical path), uses to fit the focus and writes the result to file. The second reads this file and plots the data using matplotlib.

Last night rts2af failed to fit and set the focuser far off the correct position. Now includes the chi2 in its decision and if it is too large it returns the weighted mean of the distribution as focus position. This position is only a good estimation.

The first image after a day's work.
The participants of the workshop.

Initial installation and RTS2 operation of the apache1 telescope at Astronomical Observatory of the Autonomous Region of the Aosta Valley (OAVdA) (2011-05-20)

After the introduction given by Petr and Markus and Petr's modification of the telescope dirver and the addition of the device driver for the Microfocuser the observatory was ready for the first observation in the evening. At arround 9 o'clock the sky was partly clear and apache1 started making flats and entered state night at about 21:30. After rough adjusting the secondary mirror focuser the effective focus was achieved interactively by rts2-xfocusc. A first trial of rts2af autofocuser package didn't provide a focus because the current position was far away from the real minimum, but anyway it completed. The gallery of the workshop including the first RTS2 move movie of apache1 telescope can be downloaded here (all images).

First complete rts2af autofocus cycle (2011-05-10)

An FWHM of a non saturated random star on image taken right after the focus run has been completed.
The FWHM and the FLUX fit are in very good agreement (3496, 3490). A meaningful estimator for FLUX error is still under development.

RTS2 provides a hook for analyzing images right after they have been stored. The newly installed performs now two tasks in parallel:

If detects a FWHM above a given threshold it triggers a focus run executing meaning queueing target number 5 named OnTargetFocus in the queue next of service EXEC. EXEC then executes the associated script (see postgres database table scripts), which acquires a series of preconfigured images and hands them over to as soon as they are stored on the file system. As soon as all images arrived rts2af_fit_focus.C fits the FWHM and FLUX and the result is fed back to which sets the focuser device variable FOC_DEF.

The after exposre image analysis started this morning at 05:00:40 and the focuser was set within about six minutes at 05:06:59. The first image taken right after the focus run completed, confirms the FWHM of about 2.5 pixels (3 arcsec).

A typical CCD exposure for the clear optical path is in the order of 5 seconds, a read out of the 2x2 binned 2000x2000 windowed frame takes 8 seconds and the focuser has a speed of 100 tick/second. So about 3 minutes of the ellapsed 6 minutes are used to perform the whole analysis of the 13 images.

The rts2af package is not Obs. Vermes specific. It can be used by all observatories which use RTS2. In addition, this is the place where the classes are defined, operates solely on FITS images. So it can be reused in other projects as well.

The motivation behind having a good focus is twofold. There is a large temperature induced focus shift from the beginning to the end of the night which yields a FWHM as bad as 10 pixels. Having a refractor with a diameter of "only" 0.18 meter the photons should be put in as few pixels as possible.

Error bars included.

rts2af (2011-05-07)

The current release includes now error bars (standard deviation) for the FWHM data. During the past week the configuration has been tweaked and now a single focus run is done in about four minutes. It is still an open question how to treat flux. Sextractor's output FLUX_APER/FLUXERR_APER might be a better approach.

Markus Wildi, Prof. Dr. Hans Balsiger, Dr. Rolf Bütikofer , Prof. Dr. Erwin O. Flückiger (from left to right) and Lukas Zimmermann (foreground).

Visitors of High Altitude Research Stations Jungfraujoch and Gornergrat (2011-05-06))

On Friday we welcomed the guests Prof. Dr. Hans Balsiger, Prof. Dr. Erwin Flückiger and Dr. Rolf Bütikofer of the fondation High Altitude Research Stations Jungfraujoch and Gornergrat at our observatory. We made a presentation about the construction and the present state.

rts2af Python autofocus package released (2011-04-24)

After thorough on- and offline testing the rts2af autofocus has been released to RTS2 svn repository.

Main new features:

  • There is now a script which can be executed by rts2-executor as a regular target script.
  • The fitted focus position of the optical clear path is fed back to the focuser (configurable).
  • Each image is analysed immediately after it has been stored in the file system. A few seconds after the last file has been written, the position is available.
  • The configuration of the filters has been changed: now relative offsets are stored in the configurable array (you can always dump a configuration file like it is done in SExtractor).
  • After the focus position of the clear path has been set the filtered images are done at the correct positions given you have reasonably defined filter offsets.
  • It is fast: analysing 20 4096x4096 Fits images is done in less than 10 seconds (Intel(R) Core(TM)2 Duo CPU E7400 @ 2.80GHz).
Fitted data from the focus run. The FWHM at the minimum is about 2 pixels (2.8 arcsec) for the 39 selected objects. The fitted flux peak is very close to the fwhm minimum indicating a good seeing.

Focus - Fit loop closed (2011-04-22)

After several attempts this week finally the based scripts managed to fit and set the focus.

2011-04-22T23:42:33.042 CET EXEC 4 rts2af_acquire: got fitted focuser position at: 3650 
2011-04-22T23:42:33.042 CET FOC_FLI 4 changing focuser position to 4450.000000
2011-04-22T23:42:33.611 CET FOC_FLI 4 focuser moved to 4450.000000
2011-04-22T23:42:37.045 CET FOC_FLI 4 changing focuser position to 3650.000000
2011-04-22T23:42:45.090 CET FOC_FLI 4 focuser moved to 3650.000000
2011-04-22T23:42:57.065 CET EXEC 4 rts2af_acquire: current foc_pos: 3650.0

Basically is executed by EXEC as target 3, and it spawns as a subprocess for each filter in use. The analysis subprocess reads the fits file names on its standard input and processes each file immediately. So no time is wasted. As soon as the last file appeared the fitting is done with rts2_fit_focus.C which is built on CERN's root package. Having now experience with several 50 focus runs rts2_fit_focus.C proved to be a reliable component.

Seminar on Autonomous Observatories (2011-04-17)

Markus Wildi will give a seminar on autonomous observatories at Astronomical Observatory of the Autonomous Region of the Aosta Valley (OAVdA) on May 20 through 23.

Some times one has simply to trust the results. Settings for the image on the left.
NGC 3718 and 3729 with astrometrical calibration (seen in DS9).

Cirrus clouds (2011-04-02)

Last night there were high clouds in the first half of the night. Usually the image background intensity is about 3200 units, while on the image on the left it is around 4000. The reason for the 4 bad images: one has no stars on it and in case of the other three the mount did not track.

The reason for the pronounced asymmetry of the FWHM has to be investigated. Nevertheless the fitted minimum represented the focus last night.

From 2011-04-01T01:00:00.000 CET to 2011-04-03T13:00:00.000 CET
Number of nights 2 observations nights 2 (100.00%)
Images 107 good images 104 ( 97.20%)
 Date                          Images      % of good
                    Observations   Good Images
 2011-04-01T01:00:00.000 CET    13    42    42 100.00
 2011-04-02T01:00:00.000 CET    14    65    62  95.38

Six nights in a row (2011-03-26)

The past six nights were clear and the observatory worked with the exception of Monday seamless until it closed the door in the morning. The ratio of good images varied between 2, at the beginning of the week, and 97%. The main source of the losses were incomplete focus settings in older target scripts.

Over 900 images were acquired with exposure times of typically 300 seconds. In addition 380 darks and 350 flats were exposed during twilight. The FWHM of the stars on the images was on the average 4 pixels which is about twice as much when it is really focussed.

miyo:/usr/local/src/rts-2-head # rts2-nightreport -S -f 2011-03-20 -t 2011-03-27
From 2011-03-20T00:00:00.000 CET to 2011-03-27T00:00:00.000 CET
Number of nights 7 observations nights 7 (100.00%)
Images 908 good images 470 ( 51.76%)                
                                           Good Images
                                                 % of good
 Date                 Observations
 2011-03-20T00:00:00.000 CET    62   105     4   3.81
 2011-03-21T00:00:00.000 CET    34   213     5   2.35
 2011-03-22T00:00:00.000 CET    21    83    58  69.88
 2011-03-23T00:00:00.000 CET    31   109   103  94.50
 2011-03-24T00:00:00.000 CET    45   195   104  53.33
 2011-03-25T00:00:00.000 CET    39   133   128  96.24
 2011-03-26T00:00:00.000 CET    16    70    68  97.14

There are still minor issues with the overhauled, thanks to Petr it has now a separate configuration, not sending an email report in the evening. During the past weeks the troubles with the newly introduced enforcement of the BLOCK EXPOSURE could be sorted out and rts2-dome-door-vermes as well as rts2-cupola-vermes signal their states now correctly.


Power outage during observation (2011-03-06)

In the night of 2011-03-04 we had a power outage lasting for about 15 minutes.

After the analysis of the log file it is sure that the door closed and after the failiure was over the observation continued.

Five nights in a row (2011-02-11)

M31, filters B, V, R, Halpha, 5 minutes each. Images R and Halpha were added into the red channel.

After a long period awful weather and a lot of snow we had five clear nights in row. The image of M31 on the right side was taken with the focus settings of the previous night. The average of images having an astrometrical calibration (using was over 90% indicating a smooth operation of all soft- and hardware components.The processed image with the full resolution can be found here as well as a version without Halpha filter.

The web server of the observatory has been updated and displays now the images of two dome web cams. In case you see nothing turn on the light. Be sure not doing that during clear nights.
The first production release of and were released on 2011-01-22.